OpenStudy (anonymous):

Consecutive integers are integers that follow each other in order (for example 5, 6, and 7). The sum of three consecutive integers is 417. Let n be the first one. Write an equation that will determine the three integers.

hartnn (hartnn):

first number =n so, next number = n+1 and third is n+2 sum = n+ (n+1)+ (n+2) and its given to be 417 can u form an equation now ?

OpenStudy (anonymous):

We know the numbers to equal 417 we just need to write an equation

hartnn (hartnn):

i know, its given that sum =417 and sum is n+ (n+1)+ (n+2) so the equation will be ?

OpenStudy (anonymous):

n+(n+1)+(n+2)

OpenStudy (anonymous):

?

hartnn (hartnn):

the equation will be n+n+1+n+2 = 417 so, 3n + 3=417 got this ?

hartnn (hartnn):

no ?

OpenStudy (anonymous):

Thanks

hartnn (hartnn):

welcome ^_^